# Vienna Mosaic: Navigating Social Borders in a Melting Pot

### v1

@@ Abstract
Urban segregation poses a critical challenge in cities, exacerbating inequalities, social tensions, fears, and polarization. It emerges from a complex interplay of socioeconomic disparities and residential preferences, disproportionately impacting migrant communities. In this paper, using a comprehensive administrative data from Vienna, where nearly 40% of the population consists of international migrants, we analyse co-residence preferences between migrants and locals at the neighbourhood level. Our findings reveal two major clusters in Vienna shaped by wealth disparities, district diversity, and nationality-based homophily. These insights shed light on the underlying mechanisms of urban segregation and designing policies for better integration.
@@ Introduction
Cities offer unparalleled access to essential services, including healthcare, education, transportation, cultural amenities, water and sanitation services [1, 2, 3]. Cities are the hubs where people have increased social interactions, fostering productivity and attracting talent [4]. Some of those positive urban features are further accentuated in large cities. For example, large cities tend to pull more people, primarily young [5] and have a much bigger share of international migrants [6]. Although large cities are an ideal space for social mixing and cultural integration, they can also form highly unequal and segregated communities, where groups rarely interact with others [7]. In some cities, for example, wealthy neighbourhoods tend to segregate themselves from the urban population and build themselves private cities [8]. Segregated cities often result from the residential selection process, where even small preferences for some neighbour characteristics highly alter the distribution of population groups [9, 10, 11].
@@
The emergent phenomenon of segregation greatly impacts the living spaces and design planning of cities [12]. In principle, most cities are not designed to be segregated, but individual residential selections, governmental programs [13], mobility patterns [14], and preferences to live near similar others may result in segregated cities [9]. On the positive side, segregation has been studied as a mechanism that communities use to create and preserve social capital [15]. Fueled by trust, mutual understanding, shared values, and behaviours, the social capital coming from segregation aims for cultural preservation for immigrant and local communities and to maintain socio-economic positions [16, 17, 8]. However, high levels of economic, social, and geographical segregation accentuate social, health, and education inequalities [18]. Underprivileged communities are at risk of isolation, marginalization, and victimization [19, 20] and intergroup conflict can emerge from hostility and cultural misunderstandings [21, 22]. The interplay between those different dimensions of segregation creates the need for global cities to avoid becoming disintegrated ghettos and tackle the effect of segregation on increasing polarisation, political extremism, and perception biases towards minorities [23, 24, 25].
@@
Vienna is the capital, largest city, and one of nine federal states of Austria, with about two million inhabitants spread among its 23 districts [26]. Its central position makes Vienna a key migration hub in Europe and a unique case for studying segregation patterns, particularly due to its historical context and its distinction as a leading city for quality of life [27], achieved through policies promoting affordable housing and social integration. In 2023, nearly 40% of Austria’s migrants—defined as individuals with foreign nationalities—resided in Vienna, highlighting the city’s importance as a destination for international migrants [26]. This is especially true for refugees, who often choose to resettle in Vienna due to the city’s established support systems and services for asylum seekers [28]. Vienna also stands out for its extensive history of social housing, with one of the largest stocks of community-owned buildings in Europe [29, 30]. This combination of a high migrant concentration and a strong tradition of public housing creates a unique environment in which the dynamics of integration and segregation can be explored in depth. However, despite the city’s high level of diversity, the extent of segregation within Vienna remains unclear. A comprehensive understanding of how citizens from various backgrounds integrate is still lacking. While some studies suggest spatially segregated patterns of migration [29, 31], the whole dynamics of residential integration have yet to be fully understood.
@@ Results
To explore the boundaries of socio-spatial segregation, we analyse the rates at which individuals from different nationalities reside in the same districts. First, we map all the co-living links between countries based on data on individuals who live in the same districts. Next, using statistical tests, we extract the significance level of each co-residence link, allowing us to remove random associations due to district and population sizes. Finally, we use a network community detection method based on random walks to uncover groups of countries from pairwise connections. This approach provides an interpretable map of the large-scale structure of socio-spatial segregation [32].
@@
Using raw co-living counts introduces bias due to the varying size of populations per country and district. This leads to connections driven by large inhabitant communities rather than genuinely shared interests in co-living. To account for this, we employ a statistical validation method that filters out connections likely resulting from size effects or random placement. As illustrated in Fig. 2a, we control for these random effects by simulating random relocations of residents that keep the total populations of countries and districts fixed. We then use the randomised resident counts as a baseline. Suppose the empirical data show that individuals from two countries co-reside in the same district more frequently than expected under the randomised baseline. In that case, it indicates a positive association between the two countries, suggesting a preference or affinity for shared living spaces. Conversely, if the co-residence rate is lower than expected, it points to a negative association, implying a tendency to avoid co-living in the same districts. Formally, this method uses a multinomial distribution to provide analytical expressions for the average and standard deviation of co-living rates, thereby identifying significant links in a bipartite system with countries on one side and districts on the other (Supplementary Fig. LABEL:fig:figs5) [32]. Other methods exist to evaluate significant correlations between entities in bipartite systems [33, 34, 35]. Nonetheless, our model offers the advantage of easily accounting for size variations within districts and across country populations.
@@
The z-scores help compare weights, as they account for the significant variations present in the district’s residence patterns. We then aggregate the z-scores over all districts to obtain the cumulative z-score for a pair of countries $i$ and $j$ (see Fig. 2a). Among the 21 groups (19 most populous migrant nationalities, Others, and Austrians), we identified 210 significant links, forming a fully connected, undirected, and weighted (z-scores) network based on districts co-residence. We present the network in Fig. 2b, with 80 positive links shown in red and 130 negative links in blue (see Methods and GitHub repository [36] for a detailed list of z-score values).
@@
To determine groups of countries with similar residence preferences, we use a community-detection method known as the map equation [37, 38], and its associated search algorithm, Infomap [39] (see Methods section). Infomap uses the weights of links (z-scores) to determine the clusters of strongly connected countries (see Fig. 2c). To gain better insights into the shape of these two clusters, we examine the country network structure within each cluster (see Fig. 2d and GitHub repository [36] for a detailed Infomap network). Notably, countries with larger link strengths serve as hubs, connecting to many other countries. These robust connections act as key channels for information exchange, underscoring the significant role these countries play in the dissemination of multicultural information within each cluster. People with the same nationalities may connect based on their shared background, and people living in the same districts may interact due to spatial proximity. Various post-stratification techniques, including implementing the Bonferroni correction, were employed to verify the representativeness and robustness of the findings (see Supplementary Note LABEL:sec:sanity).
@@
The key question is what factors determine the segregation of countries into two clusters. The world map of residence preferences (Fig. 2c) suggests that cultural and geographic proximity could be a key factor in the division of countries [40]. Furthermore, the world network of residence preferences (Fig. 2d) reveals that connections are typically stronger between geographically proximate countries, further supporting this hypothesis. Countries within the same cluster tend to have stronger cultural ties, sharing similar languages, traditions, and social norms. This cultural homophily drives the preference for migrants from these countries to live in close proximity, reinforcing the observed segregation.
The five most populous nationalities in the larger cluster are Austria, Ukraine, Germany, Russia, and Hungary, while the most populous nationalities in the smaller cluster are Serbia, Turkey, Syria, Romania, and Poland. Most countries belonging to the big or small clusters are from culturally similar regions [40] . There are a few notable exceptions. For instance, within the “minority cluster”, Poland and Afghanistan stand out due to their relatively distant geographic locations compared to the other countries in the group. This results in these nations being among the least connected within their cluster, suggesting that other factors may account for their inclusion. Similarly, while China and Iran are part of the “majority cluster”, their cultural and geographic distance from the other member countries leads to weaker connections within the group. This suggests that while cultural homophily is a strong driver of cluster formation, other factors, such as economic reasons, may influence the clustering patterns.
@@
When a particular nationality exhibits a strong degree of homophily (with the majority of its members residing in the same districts), it can act as an anchor, attracting migrants from other countries with similar cultural backgrounds, forming diasporas [41]. This concentration not only strengthens the presence of the initial nationality within the area but also encourages nationalities with similar cultural backgrounds to settle nearby, creating pockets of cultural cohesion across the city. Therefore, this micro-level homophily, which captures preferences for co-living within one’s national group, can compound with the clustering effects observed across larger cultural groups. Next, we examine in greater detail how neighbourhood wealth disparities and diversity characteristics also contribute to this clustering pattern, acting as additional forces in shaping co-living preferences.
Figure 2: Clustering countries sharing similar residence interests. a Diagram of the method used to build the network of co-residence. b z-score network. Edges with $\lvert z_{ij}\rvert\leq 20$ are not shown for better readability. The width and colour transparency of the links is proportional to the z-score values, while the size of the nodes reflects the total population of each country, except for Austria. The network has been produced using Noverlap layout of Gephi software [42]. Two prominent clusters are marked with a dashed line. c World map of residence preferences. Countries that belong to the same cluster have the same colour. Countries coloured with grey lines are either subsumed within Others or for which we lack data on immigrants residing in Vienna. The larger cluster, comprising the majority of the population residing in Vienna, is shown in purple and referred to as the “majority cluster”. The smaller cluster, with fewer residents, is depicted in yellow and termed the “minority cluster”. d World network of residence preferences. The size of the nodes represents the total z-score of the clusters and countries. The links represent the connections between nodes obtained from the cluster analysis with Infomap [39]; the thicker the line, the stronger the connection.
@@
We approximate the wealth distribution at the district level by the average net yearly income of the inhabitants of the district (see the Supplementary Note LABEL:sec:incomerepresentativeness). Then, for each country, we compute the Pearson correlation coefficient between the proportion of the country’s population in a district and the district’s average net income (see Methods section for details). A high positive correlation indicates that the country’s residents are concentrated in wealthier districts, suggesting greater affluence. Conversely, a strong negative correlation implies that its residents tend to live in lower-income districts. This analysis quantifies the tendency of individuals from a specific country to reside in districts where the perceived socioeconomic status of other residents or themselves is marked by a higher or lower average wealth of inhabitants. We then rank the countries from highest to lowest Income-Population fraction correlation (Fig. 3a). For instance, Austria exhibits a strong positive correlation, indicating that its residents tend to prefer districts with higher average wealth, reflecting an alignment with higher socioeconomic status. In contrast, Bosnia and Herzegovina shows a negative correlation, suggesting that its residents are more likely to live in lower-income districts, potentially influenced by economic constraints or community ties.
@@
Figure 3: Socioeconomic and diversity factors underlying nationality clusters. a Correlation between district wealth and population fraction for each nationality. b Correlation between diversity and population fraction for each nationality. Nationalities are ranked by their correlation values, from highest to lowest for wealth-population fraction correlation in a, and from lowest to highest for diversity-population fraction correlation in b. The nationality clusters are colour-coded consistently with Fig. 2c.
@@
Diversity within a district may be a consequence of broader socioeconomic and cultural factors, such as income, educational background, migrant flows, and geographical proximity. Wealthier and more cosmopolitan districts may naturally attract a lower mix of nationalities due to economic constraints limiting where migrants can afford to live, decreasing the diversity in these areas, while less affluent districts may exhibit higher diversity, resulting in a more heterogeneous population. In this context, the grouping of nationalities into two distinct groups could be due to income inequalities and cultural differences, which are reflected in patterns of residential segregation and district diversity. These associations may introduce biases that further enhance the clustering effect, creating a feedback loop: as certain nationalities settle in particular districts, their presence may influence the perceived diversity of the area, either attracting more migrants from similar backgrounds and reinforcing homogeneity or encouraging further diversity and integration.
@@ Discussion
The relationship between district diversity and the clustering of country populations in Vienna is a complex and interdependent process, where diversity could act both as a cause and a consequence of segregation, generating a multifaceted feedback loop. On the one hand, diversity can influence the residential choices of different nationalities, with some inhabitants being attracted to more diverse areas due to their openness to multicultural environments or social adaptability. In such cases, citizens may gravitate toward districts where multiple ethnicities coexist, viewing this diversity as an opportunity for social integration and cultural exchange. Conversely, other nationalities may prefer less diverse neighbourhoods, seeking the comfort and familiarity of living among people from similar cultural, linguistic, or social backgrounds. This mechanism of self-segregation, or the tendency of individuals to self-select into places based on nationality or culture, drives the formation of more homogeneous clusters, reinforcing patterns of low diversity in certain districts. On the other hand, diversity within a district may also be a consequence of broader socioeconomic and cultural factors, such as income, educational background, migrant flows, and geographical proximity.
@@
Further work could analyse the temporal evolution of migrant segregation, understanding the shape and scale of the phenomenon of internal migration, particularly toward urbanisation and mobility flows between human settlements with a stochastic Agent-Based model [43] taking into account the new arrivals depending on the Diaspora model [44].Another direction would be to analyse other factors such as urban green areas, municipal housing, Urban Vulnerability index, and the impact of infrastructure on encouraging diversity. Understanding the underlying causes of residential segregation is crucial for developing policies that foster social cohesion and integration. Addressing both economic and diversity factors can help design interventions that promote more social and inclusive neighbourhoods.

### v2

# Quantifying urban socio-economic segregation through co-residence network reconstruction

@@ Abstract
Urban segregation poses a critical challenge in cities, exacerbating inequalities, social tensions, fears, and polarization. It emerges from a complex interplay of socio-economic disparities and residential preferences, disproportionately impacting migrant communities. In this paper, using a comprehensive administrative data from Vienna, where nearly 40% of the population consists of international migrants, we analyse co-residence preferences between migrants and locals at the neighbourhood level. Our findings reveal two major clusters in Vienna shaped by wealth disparities, district diversity, and nationality-based homophily. These insights shed light on the underlying mechanisms of urban segregation and designing policies for better integration.
@@ Introduction
Cities offer unparalleled access to essential services, including healthcare, education, transportation, cultural amenities, water and sanitation services (Glaeser and Maré,, 2001; Leon,, 2008; Winters,, 2011). Cities are the hubs where people have increased social interactions, fostering productivity and attracting talent (Keuschnigg et al.,, 2019). Some of those positive urban features are further accentuated in large cities. For example, large cities tend to pull more people, primarily young (Prieto-Curiel et al.,, 2022) and have a much bigger share of international migrants (Prieto-Curiel et al.,, 2018). Although large cities are an ideal space for social mixing and cultural integration, they can also form highly unequal and segregated communities, where groups rarely interact with others (Sousa and Nicosia,, 2022). In some cities, for example, wealthy neighbourhoods tend to segregate themselves from the urban population and build themselves private cities (Pieterse et al.,, 2018). Segregated cities often result from the residential selection process, where even small preferences for some neighbour characteristics highly alter the distribution of population groups (Clark and Fossett,, 2008; Bolt et al.,, 2008; Lee,, 2016).
@@
The emergent phenomenon of segregation greatly impacts the living spaces and design planning of cities (Bruch and Swait,, 2019). In principle, most cities are not designed to be segregated, but individual residential selections, governmental programs (Trounstine,, 2018), mobility patterns (Candipan et al.,, 2021), and preferences to live near similar others may result in segregated cities (Clark and Fossett,, 2008). On the positive side, segregation has been studied as a mechanism that communities use to create and preserve social capital (Bourdieu,, 1986). Fueled by trust, mutual understanding, shared values, and behaviours, the social capital coming from segregation aims for cultural preservation for immigrant and local communities and to maintain socio-economic positions (Portes and Zhou,, 1993; Light and Thomas,, 2019; Pieterse et al.,, 2018). However, high levels of economic, social, and geographical segregation accentuate social, health, and education inequalities (DeVerteuil,, 2009). Underprivileged communities are at risk of isolation, marginalization, and victimization (O’Flaherty and Sethi,, 2007; Desmond and Shollenberger,, 2015) and intergroup conflict can emerge from hostility and cultural misunderstandings (Pettigrew and Tropp,, 2006; Helbing et al.,, 2015). The interplay between those different dimensions of segregation creates the need for global cities to avoid becoming disintegrated ghettos and tackle the effect of segregation on increasing polarisation, political extremism, and perception biases towards minorities (Düring and Wolfram,, 2015; Motyl et al.,, 2014; Lee et al.,, 2019).
@@
Vienna is the capital, largest city, and one of nine federal states of Austria, with about two million inhabitants spread among its 23 districts (Statistik Austria,, 2024). Its central position makes Vienna a key migration hub in Europe and a unique case for studying segregation patterns, particularly due to its historical context and its distinction as a leading city for quality of life (Economist-Intelligence,, 2024), achieved through policies promoting affordable housing and social integration. In 2023, nearly 40% of Austria’s migrants—defined as individuals with foreign nationalities—resided in Vienna, highlighting the city’s importance as a destination for international migrants (Statistik Austria,, 2024). This is especially true for refugees, who often choose to resettle in Vienna due to the city’s established support systems and services for asylum seekers (Ali et al.,, 2024). Vienna also stands out for its extensive history of social housing, with one of the largest stocks of community-owned buildings in Europe (Kohlbacher and Reeger,, 2016; Social Housing Vienna,, 2024). This combination of a high migrant concentration and a strong tradition of public housing creates a unique environment in which the dynamics of integration and segregation can be explored in depth. However, despite the city’s high level of diversity, the extent of segregation within Vienna remains unclear. A comprehensive understanding of how citizens from various backgrounds integrate is still lacking. While some studies suggest spatially segregated patterns of migration (Kohlbacher and Reeger,, 2016, 2020), the whole dynamics of residential integration have yet to be fully understood.
@@ Results
To explore the boundaries of socio-spatial segregation, we analyse the rates at which individuals from different nationalities reside in the same districts. First, we map all the co-living links between countries based on data on individuals who live in the same districts. Next, using statistical tests, we extract the significance level of each co-residence link, allowing us to remove random associations due to district and population sizes. Finally, we use a network community detection method based on random walks to uncover groups of countries from pairwise connections. This approach provides an interpretable map of the large-scale structure of socio-spatial segregation (Karimi et al.,, 2015).
@@
Using raw co-living counts introduces bias due to the varying size of populations per country and district. This leads to connections driven by large inhabitant communities rather than genuinely shared interests in co-living. To account for this, we employ a statistical validation method that filters out connections likely resulting from size effects or random placement. As illustrated in Fig. 2a, we control for these random effects by simulating random relocations of residents that keep the total populations of countries and districts fixed. We then use the randomised resident counts as a baseline. Suppose the empirical data show that individuals from two countries co-reside in the same district more frequently than expected under the randomised baseline. In that case, it indicates a positive association between the two countries, suggesting a preference or affinity for shared living spaces. Conversely, if the co-residence rate is lower than expected, it points to a negative association, implying a tendency to avoid co-living in the same districts. Formally, this method uses a multinomial distribution to provide analytical expressions for the average and standard deviation of co-living rates, thereby identifying significant links in a bipartite system with countries on one side and districts on the other (Supplementary Fig. LABEL:fig:figs5) (Karimi et al.,, 2015). Other methods exist to evaluate significant correlations between entities in bipartite systems (Zweig and Kaufmann,, 2011; Tumminello et al.,, 2011; Serrano et al.,, 2009). Nonetheless, our model offers the advantage of easily accounting for size variations within districts and across country populations.
@@
The z-scores help compare weights, as they account for the significant variations present in the district’s residence patterns. We then aggregate the z-scores over all districts to obtain the cumulative z-score for a pair of countries $i$ and $j$ (see Fig. 2a). Among the 21 groups (19 most populous migrant nationalities, Others, and Austrians), we identified 210 significant links, forming a fully connected, undirected, and weighted (z-scores) network based on districts co-residence. We present the network in Fig. 2b, with 80 positive links shown in red and 130 negative links in blue (see Methods and GitHub repository (Parera,, 2024) for a detailed list of z-score values).
@@
To determine groups of countries with similar residence preferences, we use a community-detection method known as the map equation (Rosvall and Bergstrom,, 2008; Rosvall et al.,, 2009), and its associated search algorithm, Infomap (Edler et al.,, 2023) (see Methods section). Infomap uses the weights of links (z-scores) to determine the clusters of strongly connected countries (see Fig. 2c). To gain better insights into the shape of these two clusters, we examine the country network structure within each cluster (see Fig. 2d and GitHub repository (Parera,, 2024) for a detailed Infomap network). Notably, countries with larger link strengths serve as hubs, connecting to many other countries. These robust connections act as key channels for information exchange, underscoring the significant role these countries play in the dissemination of multicultural information within each cluster. People with the same nationalities may connect based on their shared background, and people living in the same districts may interact due to spatial proximity. Various post-stratification techniques, including implementing the Bonferroni correction, were employed to verify the representativeness and robustness of the findings (see Supplementary Note LABEL:sec:sanity).
@@
The key question is what factors determine the segregation of countries into two clusters. The world map of residence preferences (Fig. 2c) suggests that cultural and geographic proximity could be a key factor in the division of countries (Crang,, 2013). Furthermore, the world network of residence preferences (Fig. 2d) reveals that connections are typically stronger between geographically proximate countries, further supporting this hypothesis. Countries within the same cluster tend to have stronger cultural ties, sharing similar languages, traditions, and social norms. This cultural homophily drives the preference for migrants from these countries to live in close proximity, reinforcing the observed segregation.
The five most populous nationalities in the larger cluster are Austria, Ukraine, Germany, Russia, and Hungary, while the most populous nationalities in the smaller cluster are Serbia, Turkey, Syria, Romania, and Poland. Most countries belonging to the big or small clusters are from culturally similar regions (Crang,, 2013). There are a few notable exceptions. For instance, within the “minority cluster”, Poland and Afghanistan stand out due to their relatively distant geographic locations compared to the other countries in the group. This results in these nations being among the least connected within their cluster, suggesting that other factors may account for their inclusion. Similarly, while China and Iran are part of the “majority cluster”, their cultural and geographic distance from the other member countries leads to weaker connections within the group. This suggests that while cultural homophily is a strong driver of cluster formation, other factors, such as economic reasons, may influence the clustering patterns.
@@
When a particular nationality exhibits a strong degree of homophily (with the majority of its members residing in the same districts), it can act as an anchor, attracting migrants from other countries with similar cultural backgrounds, forming diasporas (Prieto-Curiel et al., 2024b, ). This concentration not only strengthens the presence of the initial nationality within the area but also encourages nationalities with similar cultural backgrounds to settle nearby, creating pockets of cultural cohesion across the city. Therefore, this micro-level homophily, which captures preferences for co-living within one’s national group, can compound with the clustering effects observed across larger cultural groups. Next, we examine in greater detail how neighbourhood wealth disparities and diversity characteristics also contribute to this clustering pattern, acting as additional forces in shaping co-living preferences.
Figure 2: Clustering countries sharing similar residence interests. a Diagram of the method used to build the network of co-residence. b z-score network. Edges with $\lvert z_{ij}\rvert\leq 20$ are not shown for better readability. The width and colour transparency of the links is proportional to the z-score values, while the size of the nodes reflects the total population of each country, except for Austria. The network has been produced using Noverlap layout of Gephi software (Bastian et al.,, 2009). Two prominent clusters are marked with a dashed line. c World map of residence preferences. Countries that belong to the same cluster have the same colour. Countries coloured with grey lines are either subsumed within Others or for which we lack data on immigrants residing in Vienna. The larger cluster, comprising the majority of the population residing in Vienna, is shown in purple and referred to as the “majority cluster”. The smaller cluster, with fewer residents, is depicted in yellow and termed the “minority cluster”. d World network of residence preferences. The size of the nodes represents the total z-score of the clusters and countries. The links represent the connections between nodes obtained from the cluster analysis with Infomap (Edler et al.,, 2023); the thicker the line, the stronger the connection.
@@
We approximate the wealth distribution at the district level by the average net yearly income of the inhabitants of the district (see the Supplementary Note LABEL:sec:incomerepresentativeness). Then, for each country, we compute the Pearson correlation coefficient between the proportion of the country’s population in a district and the district’s average net income (see Methods section for details). A high positive correlation indicates that the country’s residents are concentrated in wealthier districts, suggesting greater affluence. Conversely, a strong negative correlation implies that its residents tend to live in lower-income districts. This analysis quantifies the tendency of individuals from a specific country to reside in districts where the perceived socio-economic status of other residents or themselves is marked by a higher or lower average wealth of inhabitants. We then rank the countries from highest to lowest Income-Population fraction correlation (Fig. 3a). For instance, Austria exhibits a strong positive correlation, indicating that its residents tend to prefer districts with higher average wealth, reflecting an alignment with higher socio-economic status. In contrast, Bosnia and Herzegovina shows a negative correlation, suggesting that its residents are more likely to live in lower-income districts, potentially influenced by economic constraints or community ties.
@@
Figure 3: Socio-economic and diversity factors underlying nationality clusters. a Correlation between district wealth and population fraction for each nationality. b Correlation between diversity and population fraction for each nationality. Nationalities are ranked by their correlation values, from highest to lowest for wealth-population fraction correlation in a, and from lowest to highest for diversity-population fraction correlation in b. The nationality clusters are colour-coded consistently with Fig. 2c.
@@
Diversity within a district may be a consequence of broader socio-economic and cultural factors, such as income, educational background, migrant flows, and geographical proximity. Wealthier and more cosmopolitan districts may naturally attract a lower mix of nationalities due to economic constraints limiting where migrants can afford to live, decreasing the diversity in these areas, while less affluent districts may exhibit higher diversity, resulting in a more heterogeneous population. In this context, the grouping of nationalities into two distinct groups could be due to income inequalities and cultural differences, which are reflected in patterns of residential segregation and district diversity. These associations may introduce biases that further enhance the clustering effect, creating a feedback loop: as certain nationalities settle in particular districts, their presence may influence the perceived diversity of the area, either attracting more migrants from similar backgrounds and reinforcing homogeneity or encouraging further diversity and integration.
@@ Discussion
The relationship between district diversity and the clustering of country populations in Vienna is a complex and interdependent process, where diversity could act both as a cause and a consequence of segregation, generating a multifaceted feedback loop. On the one hand, diversity can influence the residential choices of different nationalities, with some inhabitants being attracted to more diverse areas due to their openness to multicultural environments or social adaptability. In such cases, citizens may gravitate toward districts where multiple ethnicities coexist, viewing this diversity as an opportunity for social integration and cultural exchange. Conversely, other nationalities may prefer less diverse neighbourhoods, seeking the comfort and familiarity of living among people from similar cultural, linguistic, or social backgrounds. This mechanism of self-segregation, or the tendency of individuals to self-select into places based on nationality or culture, drives the formation of more homogeneous clusters, reinforcing patterns of low diversity in certain districts. On the other hand, diversity within a district may also be a consequence of broader socio-economic and cultural factors, such as income, educational background, migrant flows, and geographical proximity.
@@
Further work could analyse the temporal evolution of migrant segregation, understanding the shape and scale of the phenomenon of internal migration, particularly toward urbanisation and mobility flows between human settlements with a stochastic Agent-Based model (Zuccotti et al.,, 2023) taking into account the new arrivals depending on the Diaspora model (Prieto-Curiel et al., 2024a, ). Another direction would be to analyse other factors such as urban green areas, municipal housing, Urban Vulnerability index, and the impact of infrastructure on encouraging diversity. Understanding the underlying causes of residential segregation is crucial for developing policies that foster social cohesion and integration. Addressing both economic and diversity factors can help design interventions that promote more social and inclusive neighbourhoods.